Answer questions 3 through 6 based on the following information: There are two and Foreign, in a hypothetical world. Each of these two countries are Ricardian economies. Each country is endowed with 100 labor hours. Each country and car. The following table shows labor requirements per unit of each good in each country. Perfect competition prevails everywhere in the two-country world. Suppose that w and wrepresent Homes and Foreigns hourly wage rates, respectively. countries, Home can use labor to produce two goods, cheese Product Home Car (Y) Table 1: Labor hours per unit of outputs Foreign Cheese (X) Lx-2 LX aty 0.5 aLY = 2
3. The marginal cost of cheese isIn Home A) B) 2u C) 0.3r D) 4w 4. The marginal cost of cheese , in Foreign. A) t B) 2t C) 0.5 D) 4 5. The marginal cost of car isin Home. A) u B) 2w C) O.5 D) 4w 6. The marginal cost of car is in Foreign. A) B) 2w C) 0.5u D) 4w 7. What is the opportunity cost of producing one unit of cheese in Home? A) 1 car B) 2u C) 2 cars D) w 8. What is the opportunity cost of producing one unit of cheese in Foreign? A) 1 car B) 2u C) 2 cars D) w
9 9, which of the following statements is correct? A) Home is more technologically advanced than Foreign in each in B) Home is more technologically advanced than Foreign in the car industry C) Foreign commands absolute advantage over Home in each industry D) Foreign commands absolute advantage over Home only in the car 10. Which of the following statements is correct? A) Foreign is more cost effective than Home in producing both cheese and car. B) Home is more cost effective than Foreign in producing both cheese and car. C) Home is more comparatively advantaged than Foreign in producing cheese. D) Home has no comparative advantage over Foreign in each industry 11. Home and Foreign are in autarky. If you lived in Home, how much would you need to pay for a car (good Y)? A) 0.25x B) 0.5x C) 1x D) none of the above 12. Home and Foreign are in autarky. If you lived in Foreign, how much would you need to pay for a car (good Y)? A) 0.25x B) 0.5x C) 1x D) none of the above
Answer #1

Answer : 3) The answer is option B.

Because based on given informations for home country 1 unit cheese production requires 2 labor hours. Per labor hour wage rate is w. So, in home country the marginal cost is (2*w) = 2w for cheese. Therefore, option B is correct.

4) The answer is option A.

Because in foreign country 1 unit cheese production requires 1 labor hour. Per labor hour wage rate is w*. So, in foreign country the marginal cost is (1 * w*) = w*. Therefore, option A is correct.

5) The answer is option B.

Because based on given informations for home country 1 unit car production requires 2 labor hours. Per labor hour wage rate is w. So, in home country the marginal cost is (2*w) = 2w for car. Therefore, option B is correct.

6) The answer is option C.

Because based on given informations for foreign country 1 unit car production requires 0.5 labor hours. Per labor hour wage rate is w*. So, in foreign country the marginal cost is (0.5 * w*) = 0.5w* for car. Therefore, option C is correct.

7) The answer is option A.

Because in home country both per unit cheese productin and per unit car production requires 2 labor hours separately. This means that if home country produce 1 unit cheese by using 2 labor hours then the country loses the production of 1 car. So, in home country the opportunity cost of producing 1 unit cheese  is 1 car.

8) The answer is option C.

Because in foreign country 1 unit cheese production requires 1 labor hour whereas 1 unit car production requires only 0.5 labor hour. This means that by using 1 labor hour the foreign country can produce 2 cars instead of producing cheese. Therefore, in foreign country the opportunity cost of producing 1 unit cheese is 2 cars.